OpenStudy (anonymous):

Use the function rule f(x) = 3x - 2. Find the output f(1.5).

OpenStudy (anonymous):

I'm really stuck on this; could anyone possibly walk me through on how to solve this?

ganeshie8 (ganeshie8):

f(x) = 3x - 2 f(1.5) = ?

ganeshie8 (ganeshie8):

you need to plugin 1.5 in place of x

OpenStudy (anonymous):

So 1.5 = 3x - 2?

ganeshie8 (ganeshie8):

no

OpenStudy (anonymous):

Oh wait, 3 x 1.5

OpenStudy (anonymous):

I was using the wrong x, haha.

ganeshie8 (ganeshie8):

-2 also..

ganeshie8 (ganeshie8):

f(x) = 3x - 2 f(1.5) = 3*1.5 - 2 = ?

OpenStudy (anonymous):

All together it's 2.5

ganeshie8 (ganeshie8):

perfect !

OpenStudy (anonymous):

Ah! Thank you! That was so easy!

ganeshie8 (ganeshie8):

good :) np :)